OpenStudy (anonymous):

HELP!!! Integration by Partial Fractions! 1.) integral dx/(x^2+x-2) 2.) integral x^2+x-16/(x+1)(x-3)^2 dx 3.) Integral 2x^2 - 9x - 9/x^3-9x dx

OpenStudy (anonymous):

1.) \[\int\limits \frac{ dx }{ x^2+x-2 }\] 2.) \[\int\limits \frac{ x^2+x-16 }{ (x+1)(x-3)^2 } dx\] 3.) \[\int\limits \frac{ 2x^2-9x-9 }{ x^3-9x } dx\]

OpenStudy (anonymous):

The partial fractions from Mathematica follow. You do the integrations.\[\left\{\frac{1}{3 (x-1)}-\frac{1}{3 (x+2)},-\frac{1}{x+1}+\frac{2}{x-3}-\frac{1}{(x-3)^2},\frac{1}{x}+\frac{2}{x+3}-\frac{1}{x-3}\right\} \]

zepdrix (zepdrix):

So for \(\Large 1\), do you understand how to `factor` the denominator?

OpenStudy (anonymous):

Yes I do! It's \[(x+2)(x-1)\]

zepdrix (zepdrix):

Ok good good. The idea is that, there's some process (I wouldn't be able to explain where it comes from :P) that allows us to split up the factors. The numerators will always be `one degree less than` the denominator. So in this example, both factors are `linear` (x to the first power), so the unknown numerators should be `constant` \[\Large \frac{1}{(x+2)(x-1)} \quad=\quad \frac{A}{x+2}+\frac{B}{x-1}\]

zepdrix (zepdrix):

Our next step is to multiply both sides by the big chunky denominator on the left side there. k?

OpenStudy (anonymous):

Yeah! Sorry for the delay! Open Study with technical problems again! But, so far, so good!

zepdrix (zepdrix):

So multiplying both sides by the stuff gives us,\[\Large 1 \quad=\quad A(x-1)+B(x+2)\] Understand how one of the factors cancelled on the A term and one on the B term?

OpenStudy (anonymous):

Yes, I do!

zepdrix (zepdrix):

From here there are a couple of different approaches you can take to solve for A and B. You can distribute everything out, and then equate like terms. So for this problem:\[\Large 0x+1 \quad=\quad Ax-A+Bx+2B\]So you would equate the x's,\[\Large 0x=Ax+Bx \qquad\to\qquad 0=A+B\]and equate the constants,\[\Large 1=A+2B\] And from there you have a pretty easy system of equations that allows you to solve for A and B. I don't particularly care for that method though. Since both of our factors were linear we can take a nice simple approach.

zepdrix (zepdrix):

Let's start by plugging in x=1, that should give us our B value pretty quickly.

zepdrix (zepdrix):

\[\Large 1 \quad=\quad A(1-1)+B(1+2)\]See how that gets us our B nice and cleanly? :D

zepdrix (zepdrix):

You're trying to use nice values that will zero out one of the variables.

OpenStudy (anonymous):

Okay! So, it's B equal to 1/3?

zepdrix (zepdrix):

Looks good! What x value should we use to solve for A?

OpenStudy (anonymous):

We can use 1 too, right?

zepdrix (zepdrix):

Since we've found our B, we `could` plug in B=1/3, and then reuse the x=1 again, yes. But a bit of an easier way it to use an x value that will zero out the B term.

OpenStudy (anonymous):

Oh, that's true!

zepdrix (zepdrix):

\[\Large 1 \quad=\quad A(x-1)+B(x+2)\] Hmm it looks like x=-2, yes? :o

OpenStudy (anonymous):

Yes! x=-2!

zepdrix (zepdrix):

\[\Large 1\quad=\quad A(-2-1)+\cancel{B(-2+2)}\]

OpenStudy (anonymous):

So, A = -1/3

zepdrix (zepdrix):

Cool! c:

zepdrix (zepdrix):

So we've determined that:\[\Large \int\limits \frac{1}{(x+2)(x-1)}dx \quad=\quad \int\limits\frac{A}{x+2}+\frac{B}{x-1}dx\]Plugging in our A and B,\[\Large \int\limits \frac{1}{(x+2)(x-1)}dx \quad=\quad \int\limits\frac{-1/3}{x+2}+\frac{1/3}{x-1}dx\]

zepdrix (zepdrix):

We might might to factor out the 1/3 :P might make it easier to read.\[\Large \frac{1}{3}\int\limits \frac{1}{x-1}-\frac{1}{x+2}\;dx\]

OpenStudy (anonymous):

Good!

zepdrix (zepdrix):

lol i said might like 3 times in that sentence XD silly typo...

OpenStudy (anonymous):

Hahahah! Lol

zepdrix (zepdrix):

Do you understand how to integrate it from there? What types of functions should we end up with from those two bad boys?

OpenStudy (anonymous):

So, the final answer is \[\frac{ 1 }{ 3 }\ln |x-1| - \ln |x+2|\]

OpenStudy (anonymous):

But, we can simplify that!

zepdrix (zepdrix):

Yes, you might want to put brackets to show the 1/3 is multiplying both terms though. Yes, simplify! :)

zepdrix (zepdrix):

Ugh, OpenStudy is tweaking out again -_-

OpenStudy (anonymous):

Can you please show the simplification for this answer? Is it using the Log Rule?

zepdrix (zepdrix):

Yes, log rule.\[\Large \color{royalblue}{\log(a)-\log(b) \quad=\quad \log\left(\frac{a}{b}\right)}\]

zepdrix (zepdrix):

\[\Large \frac{1}{3}\left(\ln |x-1| - \ln |x+2|\right) \quad=\quad \frac{1}{3}\ln\left|\frac{x-1}{x+2}\right|\] Sorry the site is being fussy again >:O grrr You get something like this, yes? :O

OpenStudy (anonymous):

It's fine! Don't worry! That is exactly what I got!

zepdrix (zepdrix):

You could use another rule of logs to bring the 1/3 in as an exponent. But I don't think you want to do that. It looks better like this.

OpenStudy (anonymous):

Yes, it looks fine like that!

zepdrix (zepdrix):

I have to leave for a lil bit D: soz

OpenStudy (anonymous):

It's fine! But, you will be back, right?

zepdrix (zepdrix):

ya in like 45 mins prolly

OpenStudy (anonymous):

Okay, perfect! I'll be here by that time! Thank you! Numbers 2 an 3 are really complicated!

zepdrix (zepdrix):

Yessssssssssss! I got my new keyboard in the mail today!! :) So excited!!

zepdrix (zepdrix):

http://www.duckychannel.com.tw/en/Shine_3_DK9008.html So beautiful! :OOO

OpenStudy (anonymous):

AWESOME!!!

OpenStudy (anonymous):

That is so cool! :-)

zepdrix (zepdrix):

Was super expensive :O But I've been dying to get a mechanical keyboard for a while now :3

zepdrix (zepdrix):

ok ok ok back to the math XD

OpenStudy (anonymous):

Hahahah! Okay, let's go back to math!

zepdrix (zepdrix):

\[\Large \int\limits\limits \frac{ x^2+x-16 }{ (x+1)(x-3)^2 }\;dx\]Hmm so what's going on with the numerator there? :O does that factor?

zepdrix (zepdrix):

Hmm doesn't appear to factor, right? So we'll just leave it alone I guess.

zepdrix (zepdrix):

This one is a tad trickier. We have a `repeated` linear factor in the denominator. (x+1)(x-3)(x-3). See the repeat? :O So we have to do something a little sneakier with our setup.\[\Large \frac{x^2+x-16}{(x+1)(x-3)^2} \quad=\quad \frac{A}{x+1}+\frac{B}{x-3}+\frac{C}{(x-3)^2}\]

OpenStudy (anonymous):

So far, so good! :-)

zepdrix (zepdrix):

I'm not exactly sure how to explain the setup... I never really got a good explanation as to WHY we do it like that. But you have to account for each copy of the linear factor. For example if it had been (x-3)^3, we would need ANOTHER fraction, +D/(x-3)^3.

OpenStudy (anonymous):

Yes! So far, so good! :-)

zepdrix (zepdrix):

Ok so when we multiply through by our denominator we get:\[\Large x^2+x-16 \quad=\quad A(x-3)^2+B(x+1)(x-3)+C(x+1)\]

OpenStudy (anonymous):

Okay, goo!

OpenStudy (anonymous):

good!

zepdrix (zepdrix):

gooo! there's goo everywhere! +_+

zepdrix (zepdrix):

Remember earlier how we had gone over `2 methods` we could use from this point? This problem is a really good example of one in which we REALLY want to use that second method (plugging in x values). The other method: Expanding everything and then getting a system of 3 equations, would be a ton of work.

zepdrix (zepdrix):

So let's start by plugging in x=3. Looks like that should cancel out 2 of the variables for us, yes?

OpenStudy (anonymous):

Yes!

zepdrix (zepdrix):

good good good :O so what you get for C?

OpenStudy (anonymous):

1 sec

OpenStudy (anonymous):

C = -1, right?

zepdrix (zepdrix):

ya sounds good.

zepdrix (zepdrix):

Hmm looks like we can get the A value by plugging in x=-1

OpenStudy (anonymous):

A = -1/4, right?

zepdrix (zepdrix):

Sorry was fiddling with keyboard +_+ Hmm I think A=1

zepdrix (zepdrix):

On the left, the 1 and -1 cancel out.\[\Large -16 \quad=\quad A(-1-3)^2\]

OpenStudy (anonymous):

Oh, yeah! I got it! That's right!

zepdrix (zepdrix):

So we don't have a nice value we can plug in to solve for B. So let's just plug in something simple like x=0. We'll plug in our A and C values at this point also.

zepdrix (zepdrix):

\[\Large 0^2+0-16 \quad=\quad 1(0-3)^2+B(0+1)(0-3)-1(0+1)\]

OpenStudy (anonymous):

Got it!

OpenStudy (anonymous):

B = 8, right?

zepdrix (zepdrix):

mmm yah that sounds good.

OpenStudy (anonymous):

Got it!

OpenStudy (anonymous):

Still there?

OpenStudy (anonymous):

@zepdrix Outstanding work and patience.

zepdrix (zepdrix):

:o

zepdrix (zepdrix):

So I guess now that we've found our A, B and C, we should plug them in and the integrate, yah? :D\[\large \int\limits \frac{x^2+x-16}{(x+1)(x-3)^2}dx \quad=\quad \int\limits \frac{A}{x+1}+\frac{B}{x-3}+\frac{C}{(x-3)^2}dx\]

OpenStudy (anonymous):

I got this random answer form the integration!

OpenStudy (anonymous):

I got \[-\ln |x+1|+8\ln |x-3|+\frac{ 1 }{ x-3 } + C\]

zepdrix (zepdrix):

Mmm I think you have your A and C plugged in backwards.

OpenStudy (anonymous):

Oh, so it should be \[\ln |x+1|...\]

OpenStudy (anonymous):

Wait, both A and C = -1!

zepdrix (zepdrix):

Bahhh A=-1, crap my bad. Was trying to do that step in my head :P That's what I get.

zepdrix (zepdrix):

What do you mean you got a `random thing` from integrating XD lol Looks like you got the correct answer :D

OpenStudy (anonymous):

So, the first answer I wrote is right?

zepdrix (zepdrix):

mmmmmmm ya i think so

OpenStudy (anonymous):

So, is it that fine, or can we simplify it more?

zepdrix (zepdrix):

Mmmm you can combine the 2 logs if you want I suppose. Applying a rule of logs gives us:\[\Large -\ln |x+1|+8\ln |x-3| \quad=\quad \ln\frac{1}{(x+1)}+\ln(x-3)^8\]

zepdrix (zepdrix):

\[\Large= \ln\frac{(x-3)^8}{|x+1|}\]

zepdrix (zepdrix):

But I dunno, that doesn't really seem necessary :\

zepdrix (zepdrix):

Ok too much math +_+ I need a break for a bit lol

OpenStudy (anonymous):

Okay, perfect! Don't worry! Take your time!

OpenStudy (anonymous):

You are back!